# Split and Rephrase with Large Language Models

### v2

@@ 5.1 Fine-tuned LLM Variants
We first evaluated the Pythia LLM variants fine-tuned on the selected DeSSE and BiSECT training data. The results are shown in Figure 1.
In terms of parameter size, on all three metrics there is a slight tendency towards increased accuracy as the number of parameters increases, although the differences are relatively minor overall. The sole exception are the BLEU results with models fine-tuned on BiSECT and evaluated on DeSSE. In this specific scenario, the behaviour was also more erratic overall on the SARI metric, with marked drops for the 1B model.
Fine-tuning on DeSSE data led to the highest scores on all three metrics, with an average of 84.96 for BESC, 72.63 for BLEU and 56.98 for SARI, against average scores of 74.45, 38.88 and 33.27, respectively, for models fine-tuned on BiSECT data. As previously mentioned, the DeSSE training data feature both shorter sentences and lesser domain variation, which may lead to an easier modelling of the patterns and operations to be replicated at inference time. We leave a more precise analysis of these factors for future work, as it is currently relatively unclear how to precisely determine which characteristics of the fine-tuned samples may impact controlled text generation on the selected tasks. From these results, a preliminary practical conclusion might be to favour simpler and shorter data samples to fine-tune LLMs for the SPRP task.
Domain shift was detrimental in both cases. For models tested on DeSSE, the drop from in-domain to out-of-domain fine-tuning averaged to 6.82 points on BESC, 16.53 on BLEU, and 20.67 on SARI. Although less marked, models tested on BiSECT while fine-tuned on DeSSE experienced an average drop of 2.39, 3.02 and 5.55, on BESC, BLEU and SARI, respectively. The relatively larger decrease in scores with DeSSE-tuned models may be attributed to the application of these models on BiSECT test sets, which contain longer sentences on average. However, this drop still resulted in markedly higher scores compared to the BiSECT-tuned variants on the same dataset.
@@ 5.2 Instruction-tuned Models
The zero-shot and in-context few-shot results with the Alpaca and Dolly models on the SPRP task are shown in Figure 2. The first notable result is the overall higher scores obtained with Alpaca models, on all metrics, when compared to equivalent Dolly variants. The differences were not marginal as they amounted to average differences, between comparable model sizes, of 13.06, 18.57 and 13.93 points in terms of BertScore, BLEU and SARI, respectively, on the DeSSE test set. On the BiSECT test, the average differences between comparable models amounted to 8.09, 7.23 and -0.97, with only the SARI metric leading to comparable results on this test set. The choice of a specific pretrained instruction-tuned LLM for the SPRP task thus seems of particular importance, despite their relative similarity in terms of model architecture, training objectives and dataset variety.
@@
5.3 Impact of Training Data Size
The results described in the previous sections for the best performing models were obtained with approximately 6,000 training data points. In this section, we aim to evaluate LLM accuracy on the SPRP task with either larger volumes of training data, testing the ceiling of the approach, or smaller amounts to estimate minimal volumes of training data for this method. We used the fine-tuned Pythia-6.9B models for these experiments, per the previously discussed rationale. On the BiSECT corpus, which features larger amounts of data, we performed our experiments in both directions, increasing or decreasing the size of the training datasets. For the smaller DeSSE corpus, since we had already used the complete dataset, we only explored smaller training data partitions.
The results are shown in Table 2.101010In all tables, best results are shown in bold. In both cases, scores increased across metrics as training size increased, although by smaller increments from the 3K mark upwards. Best results are split between the two largest training datasets on both BiSECT and DeSSE, indicative of a potential ceiling in terms of improvements from training data augmentation. It might also be the case that the observed plateauing resulted from a lack of variety in the added training data, but verifying this hypothesis was beyond the scope of this work.
The two domains displayed noteworthy differences. Although using all the available training data resulted in better SARI and compliance scores on DeSSE, the scores obtained with 3K were closer to the best results than to those obtained with 2K training data points. For BiSECT on the other hand, compliance is under 67% for models trained with 4K data points or less, with a significant increase to 82% when using 6K and a stabilisation above 90% with training datasets of 14K or more samples. These results are in line with the previously discussed complexity of the BiSECT datasets, which feature long sentences, machine-translated content and a variety of domains which makes the SPRP task more difficult to model, compared to the relatively simpler DeSSE datasets.
@@ 6 Comparative Evaluations
pythia-6.9B-BI BiSECT 76.12 41.30 38.93 12.92 63.59 40.27 81.55
@@ 6.1 Metrics Results
From the results in the previous sections, we selected a subset of models for comparison with the ABCD and BiSECT approaches, namely: the 6.9B Pythia models, which obtained only slightly lower results on all metrics than the 12B variant and features a comparable amount of parameters to the instruction-tuned LLMs; the Alpaca zero-shot model, representative of scenarios where no data are available for fine-tuning or in-learning; and the Alpaca few-shot model using either BiSECT and DeSSE examples. For the ABCD approach, we used both the bilinear and MLP variants described in Gao et al. (2021). The comparative results are shown in Table 3.
On both test sets, the fine-tuned Pythia models with 6.9B parameters obtained significantly better results than all alternatives. On DeSSE, the pythia-6.9B-DE model outperformed the best ABCD model (ABCD-mlp) by 10.1, 14.4, 11.8 and 5.55 points in terms of BESC, BLEU, SARI and CPL, respectively; the ABCD-bilinear model obtained the best compliance score but the lowest results on the other metrics. On the BiSECT test sets, the pythia-6.9B-BI model outperformed the BiSECT method by 3.61, 4.5 and 5.72 points on BESC, BLEU, SARI, respectively; the BiSECT model only obtained better results in terms of CPL. The LLM approach can thus provide more accurate results across the board when a split is generated, but can still lag behind alternatives in terms of enforcing SPRP hypotheses.
Although fine-tuning resulted in the best results overall, it is worth noting that instruction-tuned models prompted with only 5 in-context examples obtained competitive results across the board. Their main drawback involved the lower compliance scores, compared to all other alternatives. The fine-tuned Pythia models were thus shown to be a better alternative when training datasets of sufficient size are available, but the instruction-based models could provide a reasonable alternative. The zero-shot results were also competitive against the ABCD and BiSECT approaches, although they obtained the lowest scores overall. In the few-shot approach, one limiting factor is the maximum context size, whose exploration we left for future research.
@@ 6.2 Manual Ranking Results
For this task, we adapted the Appraise environment Federmann (2018), typically used for machine translation evaluation campaigns and asked 11 evaluators to perform two sub-tasks, each involving 100 samples randomly selected: compare model hypotheses from ABCD-mlp and Pythia-6.9B fine-tuned, on the DeSSE test set, and compare hypotheses from Pythia-6.9B fine-tuned and the BiSECT model on the BiSECT test set. Evaluators were asked to compare the hypotheses according to SPRP criteria of compliance, sensicality, grammatically, missed facts and omitted facts. They were also asked to ignore minor capitalisation and tokenisation issues, which were typically present in the output of the ABCD and BiSECT models, as detected in preliminary data inspection. [^11] See Appendix D for details on the experimental protocol. The results are shown in Table 4.
@@
Table 4: 3-way ranking results constrasting hypotheses from Pythia-6.9B (P) and ABCD or BiSECT on the DeSSE and BiSECT test sets, respectively, with Krippendorf’s $\alpha$ inter-annotator agreement.
@@
The results obtained in this comparative human evaluation are thus in line with those obtained on automated metrics overall, with marked improvements achieved with LLM-based SPRP.
@@ 7 Qualitative Evaluation
To gain further insights on the SPRP capabilities of LLMs, we performed an additional human evaluation, this time centred on assessing the validity of individual SPRP hypotheses on a set of predefined criteria. For this qualitative evaluation, we sampled 100 source sentences from Wiki-BM Zhang et al. (2020a), a manually crafted benchmark created from Wikipedia data, and generated SPRP hypotheses with the Pythia-6.9B model fine-tuned on DeSSE data. The evaluation involved 8 human evaluators, who had to answer 6 questions related to the quality of an SPRP hypothesis considering a source sentence. The questions were adapted from Zhang et al. (2020a), with minor modifications: [^12] See Appendix E for details on the experimental protocol.
  1. 1.
    Is it sensical? (scale of 0-2)
  2. 2.
    Is it grammatical? (scale of 0-2)
  3. 3.
    Does it miss any existing facts? (yes/no)
  4. 4.
    Does it introduce new facts? (yes/no)
  5. 5.
    Does it have splits in the wrong place or no split at all? (yes/no)
  6. 6.
    Should some of its sentences be further split? (yes/no)
Overall, the results, shown in Table 5, were largely positive. Most output was judged sensical, grammatical, with almost no introduction of new facts, and correct splits in over 91% of the cases. The last three were judged in almost perfect inter-annotator agreement, with sensicality slightly below but still above 80% agreement among raters.
Although still featuring scores above 80%, the remaining aspects led to comparatively lower results and inter-annotator agreement, with missed facts and not enough splits in 15.87% and 16.75% of the cases, respectively, and inter-rater agreement between 0.55 and 0.60. Missed facts in particular were somewhat more difficult to judge, according to post-mortem feedback, as some rephrasing cases involved implicit rather than explicit logical connection between sentences, which could be rated more or less strictly. Similarly, the need for additional splits can be viewed as more subjective than other categories such as incorrect splits, as evaluators could view some grammatically correct additional splits as cases of over-splitting. [^13] Examples of LLM output from this evaluation are shown in Appendix B; metrics results on the complete Wiki-BM benchmark are indicated in Appendix C.
It is worth noting that these results were obtained with a model fine-tuned on data from a different domain than Wiki-BM, namely DeSSE, showing that knowledge extracted from one dataset can be effectively transferred to other domains.
@@
Sensical 86.12% / 1.84 0.803
Grammatical 92.62% / 1.91 0.907
No missed facts 84.13% 0.597
No new facts 99.5% 0.996
Correct split 91.13% 0.797
Enough split 83.25% 0.548
Table 5: Qualitative evaluation results with Pythia-6.9B ($\alpha$ = Krippendorf’s alpha inter-rater agreement). Answers for the first two questions are the percentage of 2 and the average score. For the other questions, we report the percentage of yes.
@@ 8 Conclusions
In this work, we evaluated a novel approach to the Split and Rephrase task, based on large language models. This approach was tested on two SPRP datasets, DeSSE and BiSECT, improving over the state of the art by significant margins on all metrics. The results of a manual 3-way ranking evaluation further confirmed human preference for LLM-based hypotheses, and a qualitative evaluation on Wiki-BM data further demonstrated the quality of the generated SPRP output.
Of the LLM variants we explored, the fine-tuned models obtained the best results overall, although it is worth noting though that the instruction-tuned LLMs outperformed or matched the state of the art using only 5 in-context examples, on all metrics but compliance. They might thus be a reasonable alternative using hand-crafted examples in languages where training corpora are lacking for the task.
Increased data size for fine-tuned models resulted in gradual improvement across metrics, although results stabilised with 6K on BiSECT and 2K on DeSSE. Increasing parameter size resulted in only minor improvements and models with approximately 7B parameters provided the more balanced results overall on the SPRP task.
@@ 9 Limitations
The first limitation of our approach is the need for pretrained LLMs for the specific languages in which the SPRP task needs to be performed. Note however that even English-centric LLMs can exhibit multilingual capability Armengol-Estapé et al. (2022), thus mitigating the issue. Additionally, multilingual models such as BLOOM Scao et al. (2022) can provide support for the task; we provide preliminary results in this sense in Appendix F. A more in-depth evaluation of LLM performance on the SPRP task, for languages with limited coverage, would be required to further assess the limitations of the approach.
A second limitation concerns the dependency on training data to fine-tune pretrained LLMs. We established that between 2K and 3K pairs of complex and split sentences might be necessary for the task, with high quality results obtained using between 4K and 6K training pairs, depending on the domain. Although this type of training data can be collected from either parallel multilingual datasets, as in BiSECT, or specific types of essays, as in DeSSE, this still requires dedicated work for new languages. Machine translating existing SPRP corpora with high quality models could provide a basis for fine-tuning in other languages, although a precise evaluation of this approach would be needed (see Appendix F for preliminary results along these lines). We also showed that instruction-tuned pretrained models could provide quality results using only 5 in-context learning pairs, which could be rapidly crafted manually.
@@ Appendix A LLM Training and Inference
Fine-tuning was performed with LoRA Hu et al. (2022) using the following default parameters: r=8, alpha=16, dropout=0.05; targeted modules were: Query, Key and Value.
We also used default generation parameters, namely: temperature=0.1; top_p=0.75; top_k=40; num_beans=4; max_new_tokens=128.
@@
Each training process was performed on 1 A40 GPU with 40GB of RAM. As reference times, fine-tuning took 06:07:59 hours for the Pythia-6.9B model on the BiSECT 6K training dataset and 48 hours over the larger 64K dataset. Fine-tuning the Pythia 12B model took 06:07:59 hours and the smaller 410M model took 00:37:51 hours, on the 6K BiSECT dataset.

### v4

@@ 5.2 Fine-tuned LLM Variants
Figure 1 shows the results for the Pythia variants fine-tuned on DeSSE (-DE) or BiSECT (-BI).
In terms of parameter size, for in-domain scenarios, on all metrics there is a tendency towards increased accuracy as the number of parameters increases, although only marginally so and with both SARI and CPL actually incurring minor losses with the largest DeSSE-tuned 12B models. Absolute scores were lower with the BiSECT-tuned model on its own test set, which is not unexpected considering the complexity of the test data in this case. CPL and SARI for the BiSECT-tuned model were also more erratic, with a marked drop for the 1B model in particular, which could be due to the challenges posed by training on the complex BiSECT datasets for smaller models.
Domain shift results provided interesting insights regarding robustness of the variants. For the BiSECT-tuned model, switching from its in-domain test set to the DeSSE test led to higher absolute scores for all metrics but CPL. This is likely due to the lesser complexity of the latter test, although the marked CPL drop could also indicate that training on the more complex BiSECT data resulted in limited portability for the critical splitting part of the task.
For DeSSE-tuned models, switching to the more complex BiSECT test did result in absolute drops on most metrics, as expected with a significant domain change. However, they achieved comparable results on CPL despite domains switch, and also achieved comparable results to the BiSECT-tuned model in terms of BESC, BLEU and SARI, while also largely outperformed it on CPL. Modelling SPRP on shorter sentences and limited domain variation, as featured in DeSSE, thus seem to be a better option for higher portability across domains.
@@ 5.3 Instruction-tuned Models
The zero-shot and in-context few-shot results with the Alpaca and Dolly models on the SPRP task are shown in Figure 2. The first notable result is the overall higher scores obtained with Alpaca models, on all metrics, when compared to equivalent Dolly variants. The differences were not marginal as they amounted to average differences, between comparable models, of 13.06, 18.57 and 13.93 points in terms of BertScore, BLEU and SARI, respectively, on the DeSSE test set. On the BiSECT test, the average differences between comparable models amounted to 8.09, 7.23 and -0.97, with only the SARI metric leading to comparable results on this test set. The choice of a specific pretrained instruction-tuned LLM for the SPRP task thus seems of particular importance, despite their relative similarity in terms of model architecture, training objectives and dataset variety.
@@
5.4 Impact of Training Data Size
The results described in the previous sections for the best performing models were obtained with approximately 6,000 training data points. In this section, we aim to evaluate LLM accuracy on the SPRP task with either larger volumes of training data, testing the ceiling of the approach, or smaller amounts to estimate minimal volumes of training data for this method. We used the fine-tuned Pythia-6.9B models for these experiments, per the previously discussed rationale.
The results are shown in Table 3. In both cases, scores increased across metrics as training size increased, although by smaller increments from the 3K mark upwards. Best results are split between the two largest training datasets on both BiSECT and DeSSE, indicative of a potential ceiling in terms of improvements from training data augmentation. It might also be the case that the observed plateauing resulted from a lack of variety in the additional training data, but verifying this hypothesis was beyond the scope of this work.
The two domains displayed noteworthy differences. Although using all the available training data resulted in better SARI and CPL scores on DeSSE, those obtained with 3K were closer to the best results than to those obtained with lower amounts. For BiSECT, the largest increase in terms of compliance was achieved at the 6K mark, with a stabilisation above 90% with training datasets of 14K or more samples. These results are in line with the previously discussed complexity of the BiSECT datasets, compared to DeSSE. It is worth noting that no model variant achieved compliance above 95.2, irrespective of training data size.
@@ 6 Comparative Evaluations
GPT-4 Turbo DeSSE 84.59 70.90 63.69 35.24 89.76 66.09 99.20
pythia-6.9B-BI (6K) BiSECT 76.12 41.30 38.93 12.92 63.59 40.27 81.55
pythia-6.9B-BI (30k) BiSECT 76.32 42.40 45.06 16.81 65.93 52.45 95.17
@@ 6.1 Metrics Results
From the results in the previous sections, we selected a subset of models for comparison with the ABCD and BiSECT approaches, namely: the best 6.9B Pythia models in terms of CPL and SARI (Section 5.4), which obtained only slightly lower results on all metrics than the 12B variant and features a comparable amount of parameters to the instruction-tuned LLMs; the Alpaca zero-shot model, representative of scenarios where no data are available for fine-tuning or in-learning; and the Alpaca few-shot model using either BiSECT and DeSSE examples. For the ABCD approach, we used both the bilinear and MLP variants.
We also included results from GPT-4 Turbo [^9] Available at https://help.openai.com/en/articles/8555510-gpt-4-turbo. Version gpt-4-1106-preview February 2024. We used the same instruction as for our other zero-shot results., as one of the largest LLMs to date, although these results should be considered with caution: training data are unknown, with possible test data leakage, and reproducibility standards are not met, considering the proprietary nature of the model. The comparative results are shown in Table 4.
On both test sets, the fine-tuned Pythia models with 6.9B parameters outperformed all alternatives. On DeSSE, the best ABCD model (ABCD-mlp) trailed by 10.1, 14.4, 11.8 and 5.55 points in terms of BESC, BLEU, SARI and CPL, respectively; the ABCD-bilinear model obtained the best compliance score but the lowest results on the other metrics. On BiSECT, the BiSECT model was outperformed by 3.81, 5.6 and 11.85 points on BESC, BLEU and SARI, respectively; it was however the most compliant model. The LLM variants can thus provide more accurate results across the board, but can still lag behind alternatives in terms of enforcing split SPRP hypotheses.
It is worth noting that instruction-tuned models prompted with only 5 in-context examples obtained competitive results across the board, although with significantly lower compliance scores overall. The zero-shot results were also competitive against the ABCD and BiSECT approaches, although they obtained the lowest scores overall. The fine-tuned Pythia models were thus shown to be a better alternative when training datasets of sufficient size are available, but the instruction-based models could provide a reasonable alternative.
Finally, although GPT-4 Turbo achieved the highest compliance, it was outperformed by our best LoRA variants on all other metrics. After manual examination of sampled output, the main reason seems to be a stronger tendency to rephrase and extend the original content. Different prompts might mitigate this issue, an option we left aside to maintain a fair comparison between models.
@@ 6.2 Manual Ranking Results
For this task, we adapted the Appraise environment Federmann (2018), typically used for machine translation evaluation campaigns, and asked 11 evaluators to perform two sub-tasks, each involving 100 samples randomly selected: compare hypotheses from fine-tuned Pythia-6.9B models with those of ABCD-mlp and BiSECT models, on the DeSSE and BiSECT test sets, respectively. Evaluators were asked to compare the hypotheses according to SPRP criteria of compliance, sensicality, grammatically, missed facts and omitted facts. [^10] See Appendix E for details on the experimental protocol. The results are shown in Table 5.
@@
Table 5: 3-way ranking results comparing hypotheses from Pythia-6.9B-DE (P) and alternative SPRP methods (S) ABCD or BiSECT, on the DeSSE and BiSECT test sets, respectively. $\alpha$ indicates Krippendorf’s inter-annotator agreement.
@@
The results obtained in this comparative human evaluation are thus in line with those obtained on automated metrics overall, with marked improvements achieved via LLM-based SPRP.
@@ 7 Qualitative Evaluation
To gain further insights on the SPRP capabilities of LLMs, we performed an additional human evaluation, this time centred on assessing the validity of individual SPRP hypotheses on a set of predefined criteria. For this qualitative evaluation, we sampled 100 source sentences from Wiki-BM Zhang et al. (2020a), a manually crafted benchmark created from Wikipedia data, and generated SPRP hypotheses with the Pythia-6.9B model fine-tuned on DeSSE data. The evaluation involved 8 human evaluators, who had to answer 6 questions related to the quality of an SPRP hypothesis considering a source sentence, addressing sencicality, grammaticality, deletion or introduction of new facts, correct splits and insufficient splits. [^11] The questions were adapted from Zhang et al. (2020a), with minor modifications. See Appendix F for details on the experimental protocol.
Overall, the results, shown in Table 6, were largely positive. Most output was judged sensical, grammatical, with almost no introduction of new facts, and correct splits in over 91% of the cases. All were judged with almost perfect or strong inter-annotator agreement, with sensicality and correct splitting slightly below at around 80% agreement.
Although still featuring scores above 80%, the remaining aspects led to comparatively lower marks, with missed facts and not enough splits in 15.87% and 16.75% of the cases, respectively, and inter-rater agreement between 0.55 and 0.60. Missed facts in particular were somewhat more difficult to judge, according to post-mortem feedback, as some rephrasing cases involved implicit rather than explicit logical connection between sentences, which could be rated more or less strictly. Similarly, the need for additional splits can be viewed as more subjective than other categories, as evaluators could view some grammatically correct additional splits as cases of over-splitting. [^12] Examples of LLM output from this evaluation are shown in Appendix B. Metrics results on the complete Wiki-BM benchmark are indicated in Appendix D.
It is worth noting that these results were obtained with a model fine-tuned on data from a different domain than Wiki-BM, namely DeSSE, showing that knowledge extracted from one dataset could be effectively transferred to other domains.
@@
Sensical? [0-2] 86.12% / 1.84 0.803
Grammatical? [0-2] 92.62% / 1.91 0.907
No missed facts? [Y/N] 84.13% 0.597
No new facts? [Y/N] 99.50% 0.996
Correct split? [Y/N] 91.13% 0.797
Enough split? [Y/N] 83.25% 0.548
Table 6: Qualitative evaluation results with Pythia-6.9B on Wiki-BM ($\alpha$ = Krippendorf’s alpha inter-rater agreement). Answers for the first two questions are the percentage of 2 and the average score on a [0-2] scale. For the other questions, we report the percentage of Yes.
@@ 8 Conclusions
In this work, we evaluated the performance of large language models on the Split and Rephrase task, improving over the state of the art by significant margins on all metrics over SPRP datasets of varying complexity. The results of a manual 3-way ranking evaluation further confirmed human preference for LLM-based hypotheses, and a qualitative evaluation on Wiki-BM data further demonstrated the quality of the generated SPRP output.
Among LLM variants, models fine-tuned with LoRA achieved the best results overall. Zero-short and few-shot instruction-tuned variants achieved reasonable performance, modulo a significant drop in splitting compliance for the latter, and might thus constitute a reasonable alternative in languages where training corpora are lacking for the task.
Increasing fine-tuning data size resulted in only gradual improvements, plateauing on most metrics with subsets of the available data. Increased parameters led to only minor improvements, with models around 7B achieving the more balanced results. Despite significant improvements, none of the examined variants achieved full task compliance, and core aspects of SPRP remain a challenge.
@@ 9 Limitations
The first limitation of an LLM-based approach is the need for pretrained LLMs for the specific languages in which the SPRP task needs to be performed. Note however that even English-centric LLMs can exhibit multilingual capability Armengol-Estapé et al. (2022), thus mitigating the issue. Additionally, multilingual models such as BLOOM Scao et al. (2022) can provide support for the task; we provide preliminary results in this sense in Appendix G. A more in-depth evaluation of LLM performance on the SPRP task, for languages with limited coverage, would be required to further assess the limitations of the approach.
A second limitation concerns the dependency on training data to fine-tune pretrained LLMs. We established that around 3K pairs of complex and split sentences might be necessary for the task, with high quality results obtained using between 4K and 6K training pairs, depending on the domain. Although this type of training data can be collected from either parallel multilingual datasets, as in BiSECT, or specific types of essays, as in DeSSE, this still requires dedicated work for new languages. Machine translating existing SPRP corpora with high quality models could provide a basis for fine-tuning in other languages, although a precise evaluation of this approach would be needed (see Appendix G for preliminary results along these lines). We also showed that instruction-tuned pretrained models could provide quality results using only 5 in-context learning pairs, which could be rapidly crafted manually.
@@ 10 Ethical Considerations
Acknowledgements
We wish to thank the anonymous ARR reviewers for their helpful comments, and the participants in our manual evaluations for their time and contribution. This work was partially supported by the Department of Economic Development and Competitiveness of the Basque Government (Spri Group) through funding for the IRAZ project (ZL-2024/00570).
@@ Appendix A LLM Training and Inference
Fine-tuning was performed with LoRA Hu et al. (2022) using the following default parameters: r=8, alpha=16, dropout=0.05; targeted modules were: Query, Key and Value. We also used default generation parameters, namely: temperature=0.1; top_p=0.75; top_k=40; num_beams=4; max_new_tokens=128.
@@
Each training process was performed on 1 A40 GPU with 48GB of RAM. As reference times, fine-tuning took 06:07:59 hours for the Pythia-6.9B model on the BiSECT 6K training dataset and 48 hours over the larger 64K dataset. Fine-tuning the Pythia 12B model took 06:07:59 hours and the smaller 410M model took 00:37:51 hours, on the 6K BiSECT dataset.
For soft-prompting methods, we used the following default settings: num_virtual_tokens=20, prompt_tuning_init="Split the sentence below into separate sentences." for Prompt-tuning and num_virtual_tokens=20, encoder_hidden_size=128 for P-tuning.